# TabularARGN: A Flexible and Efficient Auto-Regressive Framework for Generating High-Fidelity Synthetic Data

### v1

@@ Abstract
Synthetic data generation for tabular datasets must balance fidelity, efficiency, and versatility to meet the demands of real-world applications. We introduce the Tabular Auto-Regressive Generative Network (TabularARGN), a flexible framework designed to handle mixed-type, multivariate, and sequential datasets. By training on all possible conditional probabilities, TabularARGN supports advanced features such as fairness-aware generation, imputation, and conditional generation on any subset of columns. The framework achieves state-of-the-art synthetic data quality while significantly reducing training and inference times, making it ideal for large-scale datasets with diverse structures. Evaluated across established benchmarks, including realistic datasets with complex relationships, TabularARGN demonstrates its capability to synthesize high-quality data efficiently. By unifying flexibility and performance, this framework paves the way for practical synthetic data generation across industries.
Machine Learning, ICML
@@ 1 Introduction
In this paper, we revisit an existing approach for density estimation: auto-regressive models, which approximate the full joint density $p(\mathbf{x})$ as a product of conditional probabilities. This decomposition leverages the chain rule of probability, where $p(\mathbf{x})=\prod_{i=1}^{D}p(x_{i}\mid x_{1},\ldots,x_{i-1})$, breaking a complex density estimation problem into a series of simpler conditional distributions. Auto-regressive models have proven remarkably successful in density estimation tasks in natural language processing (Brown et al., 2020; OpenAI et al., 2024) and image generation (Tian et al., 2024; Sun et al., 2024), underpinning recent advances in synthetic image creation and, more prominently, the transformative success of large language models (LLMs).
While there is a growing body of literature leveraging LLM-type models (Hegselmann et al., 2023; Smolyak et al., 2024; Miletic & Sariyar, 2024), both token-based custom-built transformer architectures and fine-tuned foundational language models, for tabular synthetic data generation, these approaches often treat rows of tabular data as sequences of text. In contrast, the literature on specialized auto-regressive models designed specifically for tabular data, in its various forms and structures, remains scarce. These specialized models do not rely on treating rows as text but instead exploit the constrained value ranges of structured tabular datasets. Motivated by this gap, we introduce the Tabular Auto-Regressive Generative Network (TabularARGN) framework, which applies auto-regressiveness across the discretized column, time, and table dimensions to generate high-fidelity synthetic data for flat and sequential tables.
For the purpose of this paper, we define flat tables as datasets where each row represents an independent and identically distributed (i.i.d.) sample from an underlying distribution. In the context of privacy protection, each data subject, whose privacy is to be safeguarded, is represented by a single row. Sequential tables, by contrast, contain multiple records per data subject, grouped by a unique key provided within the tabular data. Furthermore, our presented framework supports the inclusion of additional context for each training sample through a separate table associated via a primary–foreign key one-to-many relationship. For example, a flat table could include time-independent covariates, such as a bank customer’s date of birth or membership status, while the sequential table captures the customer’s transaction history.
@@
2 Related work and our contribution
To the best of our knowledge, Uria et al. (2016) were the first to develop an auto-regressive neural network (NADE) for tackling the problem of unsupervised distribution and density estimation. This foundational idea and architecture serve as the backbone of the TabularARGN flat and sequential models, leveraging the same core principle of decomposing the joint density of a dataset into a product of conditional probabilities. The strong performance of NADE in density estimation directly translates to the strong performance of TabularARGN in generating high-quality synthetic data.
@@ 2.1 Flat Models
In their study, Uria et al. (2016) calculate the log-likelihood performance of NADE on binary tabular and binary image datasets, as well as purely real-valued datasets. However, they do not extend their work to generate synthetic data samples. Furthermore, the NADE architecture is not adapted in their work to handle (a) multi-categorical datasets with polytomous variables (categorical variables with cardinality $\geq 2$), or (b) mixed-type datasets that combine categorical and numeric variables, both of which are critical for real-world tabular data synthesis tasks.
Bayesian Networks (BNs) can be considered the predecessors of NADE, as they also model the joint distribution of a dataset by decomposing it into conditional probabilities. However, unlike NADE, BNs rely on a directed acyclic graph to represent dependencies between variables, where each node corresponds to a variable, and the edges represent conditional dependencies. This graphical structure allows BNs to explicitly encode the probabilistic relationships in a dataset. Unlike NADE, which was originally designed for density estimation, Bayesian Networks have been implemented to generate tabular synthetic data (Ankan & Panda, 2015; Qian et al., 2023)
@@
With the rise in popularity of Large Language Models (LLMs), token-based transformers have also been applied to generate tabular synthetic data. Broadly, two classes of tabular-data LLMs can be identified. The first class leverages the knowledge embedded in pre-trained foundational models to create synthetic data ”from scratch”, relying solely on user prompts. The second class focuses on generating synthetic data that replicates a specific original dataset. This approach involves either fine-tuning pre-trained LLMs on the target dataset or training LLM-like architectures from scratch using the dataset. In this paper, we benchmark only against methods following the latter approach, where the model is explicitly trained to reproduce the statistical properties of a given dataset (Borisov et al., 2023; Solatorio & Dupriez, 2023; Karami et al., 2024).
@@
Diffusion models, originally developed for image and audio synthesis, have also been adapted for generating tabular synthetic data (Lee et al., 2023; Kim et al., 2023; Kotelnikov et al., 2023; Zhang et al., 2024b; villaizánvallelado2024diffusionmodelstabulardata). These models leverage iterative denoising processes to approximate complex data distributions and have shown promise in capturing intricate patterns in tabular datasets. It has also been shown that diffusion models can effectively memorize the training data with a larger number of training epochs (Fang et al., 2024).
@@ 2.3 Our Contributions
  • Compute Efficiency: The training and sampling of TabularARGN models are orders of magnitude faster than other generative SOTA models for tabular data, making it possible to apply these to significantly larger real-world datasets.
  • Sampling Flexibility: TabularARGN supports advanced sampling capabilities, including:
    • Conditional generation on arbitrary subsets of mixed-type attributes.
    • Imputation of missing values.
    • Incorporation of statistical fairness across sensitive attributes (Krchova et al., 2023).
    • Controlling sampling probabilities via temperature adjustments to balance rule-adherence with data diversity.
  • Data Versatility: TabularARGN accommodates the heterogeneity of real-world tabular datasets, including:
    • Multi-variate, mixed-type data (categorical, numerical, date-time, geo-spatial).
    • Multi-sequence datasets with varying sequence lengths and varying time intervals.
    • Missing values.
@@
  • Open-Source Framework: We provide a well-tested and well-maintained reference implementation of TabularARGN under a fully permissive Apache v2 Open Source license at https://github.com/mostly-ai/mostlyai-engine/.
@@ 3.3 Model Components
The embedding vectors $\mathbf{e}_{x_{i}}$ for all sub-columns are then concatenated into a single vector $\left[\mathbf{e}_{x_{1}},...,\mathbf{e}_{x_{N}}\right]$, which is then passed to the permutation masking layer. The permutation layer introduces a ”causal” feeding mechanism based on a random column order - varying from batch to batch - ensuring that each column’s regressor receives only the inputs it is allowed to access based on the conditional dependencies. The masking layer modifies the concatenated vector by ”blanking out” (i.e., setting to zero) the contributions not permitted for specific regressor layers.
@@
The regressor block consists of one or more feed-forward layers (see appendix B for details), each using a ReLU activation function. To enhance generalization and prevent overfitting, a dropout is applied during training to each layer within the regressor block. The predictor layer is a single feed-forward layer with a softmax activation applied during generation, outputting normalized, discrete probability vectors that represent the estimated probabilities for each category $x_{i}^{(k)}$. The size of the predictor layer corresponds to the number of categories within the specific sub-column.
@@ 3.4 Model Training
Table Type Dataset Model Accuracy DCR share Training Time
flat adult TabularARGN 97.9% 50.3% 138 s
flat adult TabularARGN, $\epsilon=2.8$ 93.8% 50.4% 537 s
flat adult TabSyn 98.2% 50.0% 2,316 s
flat adult STaSy 80.6% 50.0% 4,588 s
flat adult CTGAN 79.0% 49.5% 2,684 s
flat acs-income TabularARGN 98.5% 50.2% 3,511 s
flat acs-income TabularARGN, $\epsilon=2.3$ 96.7% 52.5% 9,050 s
flat acs-income TabSyn 97.5% 49.7% 8,942 s
flat acs-income STaSy - - -
flat acs-income CTGAN - - -
sequential baseball TabularARGN 88.4% 50.5% 174 s
sequential baseball TabularARGN, $\epsilon=3.6$ 79.0% 51.7% 2,558 s
sequential baseball REaLTabFormer 78.0% 49.2% 86,031 s
sequential baseball RC-TGAN 70.0% 47.2% 40,564 s
sequential baseball ClavaDDPM 79.4% 49.4% 3,997 s
sequential california TabularARGN 98.1% 51.1% 6,887 s
sequential california TabularARGN, $\epsilon=1.2$ 93.4% 51.2% 18,690 s
sequential california REaLTabFormer 88.6% 49.0% 32,912 s
sequential california RC-TGAN 73.3% 47.0% 52,051 s
sequential california ClavaDDPM 77.3% 50.0% 2,597 s
Table 1: Overview of empirical results. We excluded results of STaSy in the ACS-Income data set because, despite observing a similarly smooth loss behavior as for the Adult dataset, the synthetic data quality was unexpectedly poor, and we were unable to identify the underlying cause with confidence. To avoid drawing potentially misleading conclusions, we have chosen not to include these results. Results for CT-GAN are excluded because training did not finish within 24 hours.
@@ 3.6 Conditional Sequence Generation - Two-Table setups
Figure 3: Training timings in seconds across data sets and models. Refer to the caption of Tab. 1 for an explanation of why the results for STaSy and CT-GAN are omitted.
@@ 4 Empirical Results
In this section, we compare the synthetic data quality as well as the compute efficiency of TabularARGN to available state-of-the-art (SOTA) models, across a variety of small and medium-scaled datasets. Analogous to the previous section, we will also split this section into a study on flat tables and one for sequential tables. For each, we utilize selected methods and data sets from existing benchmarks in the literature (Zhang et al., 2024b; Pang et al., 2024). These benchmarks are extended by considering a larger-scale, more realistic dataset for flat and sequential data. By ”realistic” we refer to datasets with higher dimensionality, specifically in terms of the number of rows, columns, and categories. This approach highlights the advantages of TabularARGN, particularly in its ability to efficiently handle the challenges of synthesizing large, complex datasets typical of practical applications.
@@ 4.1 Flat Tables
As baseline models, we use CTGAN (Xu et al., 2019) as implemented in (Qian et al., 2023) and STaSy (Kim et al., 2023), and TabSyn as implemented in the comprehensive TabSyn benchmark by (Zhang et al., 2024b). CTGAN was one of the first synthetic data generators and is included for reference. The other three are the best-performing methods in terms of the quality/efficiency trade-off in the TabSyn benchmark. While there are additional methods included in the TabSyn benchmark, these are not directly evaluated in our study. However, as we use a dataset that is common to both TabSyn and its recent extension TabDar (Zhang et al., 2024a), our results enable a qualitative comparison with those of these additional methods.
We evaluate TabularARGN against baseline methods on two data sets: Adult (Dua & Graff, 2019) and ACS-Income (Ding et al., 2021; Flood et al., 2020). The Adult data set has around 48k rows with eight categorical and six numerical features. ACS-Income has about 1.5 million rows, containing 28 categorical and four numerical features—making it roughly 30 times larger and twice as wide as Adult, and thus a challenging real-world benchmark. We initially considered an even larger data set with more columns, but all baseline methods were either too slow or required excessive GPU memory. Appendix F details the computing setup, benchmark methods, and test data sets.
On both datasets, TabularARGN performs on par with the state-of-the-art method TabSyn, trailing it slightly on Adult but surpassing it on ACS-Income (see Table 1). Both methods exceed 95% accuracy, clearly outperforming STaSy and CTGAN where available.
For both datasets, TabularARGN has by far the shortest training time, being faster by more than an order of magnitude compared to all other methods on Adult (see Fig.3). For ACS-Income, TabularARGN remains highly efficient, being more than twice as fast as TabSyn.
DCR shares for all methods remain near 0.5 (=50%), indicating no significant information leakage from individual training samples to the synthetic data when compared to holdout samples.
We also report results for TabularARGN runs with Differential Privacy. For Adult and ACS-Income, TabularARGN achieves competitive accuracy scores above 90% with $\epsilon=2.8,\delta=1e^{-5}$ and $\epsilon=2.3,\delta=1e^{-5}$, respectively — values corresponding to the privacy levels reached at model convergence. These results still clearly outperform STaSy and CTGAN. These strict mathematical privacy guarantees come at the cost of increased training times, with factors of $2.6$ and $3.9$ for Adult and ACS-Income, respectively.
@@ 4.2 Sequential Tables
Establishing a comprehensive benchmark for sequential tables presents challenges, as many methods proposed in the literature for sequence generation are constrained by the properties of the data they can handle. As already mentioned earlier, common limitations include the inability to process discrete features, reliance on single-sequence settings with sequences chunked during training, restrictions to constant sequence lengths across samples, and reliance on constant or equidistant time steps. These limitations make it difficult for such methods to handle real-world datasets, which often exhibit diverse structures, varying sequence lengths, and a mix of feature types.
The baseline methods selected for our study address all the aforementioned limitations - at least after some basic pre-processing - and are capable of producing synthetic data of competitive quality. We evaluate REalTabFormer (Solatorio & Dupriez, 2023), RC-TGAN (Gueye et al., 2023), and ClavaDDPM (Pang et al., 2024) which represent a diverse cross-section of generative approaches.
We compare TabularARGN with the baseline models on two data sets: a two-table subset (players and fielding) of the Baseball dataset (Lahman, 2023) and the California dataset (Pace & Barry, 1997) consisting of a table of households and a table of individuals. For Baseball, the players table (flat context) has about 20 thousand rows, and the sequential fielding table has 140 thousand, resulting in an average sequence length of seven. With over 600 thousand rows, the household table has about $30$ times more records than the players table, resulting in almost 1.7 million rows and an average sequence length of 2.8 in the individuals table. Both sequence tables have a comparable number of features: 11 (Baseball) and 15 (California). Again, more detailed information on methods and data sets can be found in appendix F.
For both sequential datasets, TabularARGN achieves the highest accuracy, surpassing all other methods by nearly 10 percentage points (see Table 1). Even with differential privacy enabled ($\epsilon=3.6,\delta=1e{-5}$ for Baseball and $\epsilon=1.2,\delta=1e{-5}$ for California), TabularARGN remains on par with the top-performing baseline methods. However, the added privacy guarantee increases training time by a factor of 10 for Baseball and 2.7 for California.
For the Baseball dataset, TabularARGN exhibits training times one to two orders of magnitude smaller than those of the baseline models (see Fig. 3). For the California dataset, the gap narrows significantly, with RealTabFormer and RC-TGAN being slower by factors of 4.8 and 7.5, respectively. ClavaDDPM, while 2.7 times faster than TabularARGN, sacrifices over 20 percentage points in accuracy, yielding only low-fidelity synthetic data.
Interestingly, RealTabFormer and ClavaDDPM train faster on the larger California dataset. Without further investigation, we hypothesize that this behavior is related to the longer sequence lengths in the Baseball dataset.
As with the flat tables, DCR shares vary only slightly around 0.5, indicating no major information leakage from the training samples to the synthetic samples, when compared to holdout samples.

### v2

@@ Abstract
Synthetic data generation for tabular datasets must balance fidelity, efficiency, and versatility to meet the demands of real-world applications. We introduce the Tabular Auto-Regressive Generative Network (TabularARGN), a flexible framework designed to handle mixed-type, multivariate, and sequential datasets. By training on randomized subsets of conditional probabilities, TabularARGN supports advanced features such as fairness-aware generation, imputation, and conditional generation on any subset of columns. The framework achieves state-of-the-art synthetic data quality while significantly reducing training and inference times, making it ideal for large-scale datasets with diverse structures. Evaluated across established benchmarks, including realistic datasets with complex relationships, TabularARGN demonstrates its capability to synthesize high-quality data efficiently. By unifying flexibility and performance, this framework paves the way for practical synthetic data generation across industries.
Machine Learning, Synthetic Data, Data Generation, Tabular Data, Generative AI, Auto-Regressive Generative Neural Network, ICML
@@ 1 Introduction
In this paper, we revisit auto-regressive models for density estimation, which approximate the full joint density $p(\mathbf{x})$ as a product of conditional probabilities. This decomposition leverages the chain rule of probability, where $p(\mathbf{x})=\prod_{i=1}^{D}p(x_{i}\mid x_{1},\ldots,x_{i-1})$, breaking a complex density estimation problem into a series of simpler conditional distributions. Auto-regressive models have proven remarkably successful in density estimation tasks in natural language processing (Brown et al., 2020; OpenAI et al., 2024) and image generation (Tian et al., 2024; Sun et al., 2024), underpinning recent advances in synthetic image creation and, more prominently, the transformative success of large language models (LLMs).
While there is a growing body of literature leveraging LLM-type models (Hegselmann et al., 2023; Smolyak et al., 2024; Miletic & Sariyar, 2024), both token-based custom-built transformer architectures and fine-tuned foundational language models for tabular synthetic data generation, these approaches often treat rows of tabular data as sequences of text. In contrast, the literature on specialized auto-regressive models designed specifically for tabular data, in its various forms and structures, remains scarce. These specialized models do not rely on treating rows as text but instead exploit the constrained value ranges of structured tabular datasets. Motivated by this gap, we introduce the Tabular Auto-Regressive Generative Network (TabularARGN) framework, which applies auto-regressiveness across the discretized column, time, and table dimensions to generate high-fidelity synthetic data for flat and sequential tables.
We define flat tables as datasets where each row represents an independent and identically distributed (i.i.d.) sample from an underlying distribution. In the context of privacy protection, each data subject, whose privacy is to be safeguarded, is represented by a single row. Sequential tables, by contrast, contain multiple records per data subject, grouped by a unique key provided within the tabular data. Furthermore, our presented framework supports the inclusion of additional context for each training sample through a separate table associated via a primary–foreign key, one-to-many relationship. For example, a flat table could include time-independent covariates, such as a bank customer’s date of birth or membership status, while the sequential table captures the customer’s transaction history.
@@
2 Related work and contributions
To the best of our knowledge, Uria et al. (2016) were the first to develop an auto-regressive neural network (NADE) for tackling the problem of unsupervised distribution and density estimation. This foundational idea and architecture serve as the backbone of the TabularARGN flat and sequential models.
@@ 2.1 Flat Models
In their study, Uria et al. (2016) calculate the log-likelihood performance of NADE on binary tabular and binary image datasets, as well as purely real-valued datasets. However, they do not extend their work to generate synthetic data samples. Furthermore, the NADE architecture is not adapted in their work to handle multi-categorical datasets with variables of cardinality $\geq 2$, nor mixed-type datasets that combine categorical and numeric variables, both of which are critical for real-world tabular data synthesis tasks. NADE shares similarities with Bayesian Networks (BNs), as both methods factorize the joint probability into conditional distributions. However, BNs represent dependencies between variables using a directed acyclic graph. Once the conditional probability distributions are learned, data can be sampled from the model, making BNs a natural choice for generating tabular data (Ankan & Panda, 2015; Qian et al., 2023).
@@
With the rise in popularity of Large Language Models (LLMs), token-based transformers have also been applied to generate tabular synthetic data based on a specific target dataset. Some methods finetune pre-trained LLMs directly, while others train LLM-like architectures from scratch (Borisov et al., 2023; Solatorio & Dupriez, 2023; Karami et al., 2024).
@@
Diffusion models, originally developed for image and audio synthesis, have also been adapted for generating tabular synthetic data (Lee et al., 2023; Kim et al., 2023; Kotelnikov et al., 2023; Zhang et al., 2024b; Villaizán-Vallelado et al., 2024). These models leverage iterative denoising processes to approximate complex data distributions and have shown promise in capturing intricate patterns in tabular datasets. It has also been shown that diffusion models can effectively memorize the training data with a larger number of training epochs (Fang et al., 2024).
@@ 2.3 Our Contributions
  • Compute Efficiency: The training and sampling of TabularARGN models are sometimes orders of magnitude faster than other generative SOTA models for tabular data, making it possible to apply these to significantly larger real-world datasets.
@@
  • Sampling Flexibility: TabularARGN supports conditional generation on arbitrary subsets of mixed-type attributes enabling data generation tailored to specific constraints. It also facilitates missing value imputation and the incorporation of statistical fairness across sensitive attributes (Krchova et al., 2023). Additionally, sampling probabilities can be adjusted via a temperature parameter to balance rule adherence with data diversity.
  • Data Versatility: TabularARGN accommodates the heterogeneity of real-world tabular datasets, including multi-variate, mixed-type data (categorical, numerical, date-time, geospatial), multi-sequence datasets with varying sequence lengths and varying time intervals, and missing values.
@@
  • Open-Source Framework: We provide a well-tested and well-maintained implementation of TabularARGN under a fully permissive Apache v2 Open Source license at https://github.com/mostly-ai/mostlyai-engine/.
@@ 3.3 Model Components
The embedding vectors $\mathbf{e}_{x_{i}}$ for all sub-columns are concatenated into a single vector $\left[\mathbf{e}_{x_{1}},...,\mathbf{e}_{x_{N}}\right]$, which is then passed to the permutation masking layer. The permutation layer introduces a ”causal” feeding mechanism based on a random column order - varying from batch to batch - ensuring that each column’s regressor receives only the inputs it is allowed to access based on the conditional dependencies. The masking layer modifies the concatenated vector by ”blanking out” (i.e., setting to zero) the contributions not permitted for specific regressor layers.
@@
The regressor block consists of one or more feed-forward layers (see appendix B for details), each using a ReLU activation function. To enhance generalization and prevent overfitting, dropout is applied during training to each layer within the regressor block. The predictor layer is a single feed-forward layer with a softmax activation applied during generation, outputting normalized, discrete probability vectors that represent the estimated probabilities for each category $x_{i}^{(k)}$. The size of the predictor layer corresponds to the number of categories within the specific sub-column.
@@ 4 Empirical Results
In this section and Appendix G, we compare the synthetic data quality as well as the compute efficiency of TabularARGN to available state-of-the-art (SOTA) models, across a variety of small and medium-scaled datasets. For maintaining focus, we present results for one representative flat dataset and one sequential dataset here, while additional results on additional datasets - demonstrating similarly strong performance of TabularARGN - are provided in Appendix G. Analogous to the previous sections, we split this section into a study on flat and sequential tables. For each, we utilize selected methods and data sets from existing benchmarks in the literature (Zhang et al., 2024b; Pang et al., 2024).
@@
In this section, we present results in terms of runtime and synthetic data quality - the overall accuracy computed as the average across univariate, bivariate, and coherence accuracy. We report the DCR shares only in Appendix G as none of the methods displays a significant increase beyond 50%.
@@ 4.1 Flat Tables
As baseline models, we include CTGAN (Xu et al., 2019), STaSy (Kim et al., 2023), TabSyn (Zhang et al., 2024b), TabMT (Gulati & Roysdon, 2023), and Unmasking Trees (McCarter, 2024), providing a strong mix of state-of-the-art models from different generative approaches, including a GAN-based method as well as diffusion and auto-regressive models. We refer the reader to Appendix F for a more detailed description of these methods, as well as the statistics of all datasets utilized in this work. On the Adult data set (Dua & Graff, 2019) - 24k rows with eight categorical and six numerical variables - TabularARGN performs on par with state-of-the-art methods such as TabSyn in terms of accuracy while outperforming all benchmarks in terms of runtime by a large margin (see Figure 3 left column). Notably, when employing Differential Privacy ($\epsilon\sim 2.5$, $\delta=10^{-5}$), performance remains competitive, with only a modest decline in accuracy, while the runtime increases moderately but remains well below that of the baselines. The reported epsilon values correspond to the privacy levels reached at model convergence. Further results on other methods and flat table datasets can be found in Appendix G.
Figure 3: Training time (top row) and Accuracy (bottom row) for the flat Adult (left column) and sequential Baseball (right column) data set. Reported values are averages over five full training and generation runs, error bars indicate the minimum and maximum values. As UnmaskingTrees is a non-Deep Learning approach, it is run on CPU. Training of TabularARGN on the Adult with a CPU takes approximately 2.5 minutes.
@@ 4.2 Sequential Tables
Establishing a comprehensive benchmark for sequential tables presents challenges, as many methods proposed in the literature for sequence generation are constrained by the properties of the data they can handle. As already mentioned earlier, common limitations include the inability to process discrete features, reliance on single-sequence settings with sequences chunked during training, restrictions to constant sequence lengths across samples, and reliance on constant or equidistant time steps. These limitations make it difficult for such methods to handle real-world datasets, which often exhibit diverse structures, varying sequence lengths, and a mix of feature types. The baseline methods selected for our study address all these limitations - at least after some basic pre-processing - and are capable of producing synthetic data of competitive quality. We evaluate REalTabFormer (Solatorio & Dupriez, 2023), RC-TGAN (Gueye et al., 2023), and ClavaDDPM (Pang et al., 2024) which represent a diverse cross-section of generative approaches.
We compare TabularARGN with the baseline models on a two-table subset (players and fielding) of the Baseball dataset (Lahman, 2023). The players table (flat context) has about 20 thousand rows, and the sequential fielding table has 140 thousand, resulting in an average sequence length of seven. More detailed information on methods and data sets can be found in appendix F.
TabularARGN achieves the highest accuracy, surpassing all other methods by nearly 10 percentage points (see Figure 3 right column). Even with differential privacy enabled ($\epsilon\sim 4$, $\delta=10^{-5}$), TabularARGN remains on par with the top-performing baseline methods. Even though adding privacy guarantees increases the training time by a factor of 10, our method remains one to two orders of magnitude below those of the baseline models, reinforcing its scalability and practical usability for real-world applications. Further results on other methods and sequential table datasets can be found in Appendix G.